OpenStudy (anonymous):

Please Help!! Write the sum using summation notation, assuming the suggested pattern continues. -9 - 4 + 1 + 6 + ... + 66

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Can you tell me the pattern (whether I know it or not)

OpenStudy (anonymous):

n+5

OpenStudy (anonymous):

i think

OpenStudy (solomonzelman):

Yes, good. You add 5 each time.

OpenStudy (solomonzelman):

66 is the last term, right ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 66=-9 + 5(x) }\) x is `which number term is 66` in the sequence. Makes sense so far ?

OpenStudy (anonymous):

yeah, so i figure out what x is now?

OpenStudy (solomonzelman):

yes, please:)

OpenStudy (anonymous):

x=15

OpenStudy (solomonzelman):

Yes, good !

OpenStudy (solomonzelman):

So how many terms are there ?

OpenStudy (anonymous):

15?

OpenStudy (solomonzelman):

So we know what to put on top of the sigma, 15.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\sum_{n=0}^{15}\left( -9+5n \right)\] is this what it will look like?

OpenStudy (solomonzelman):

well, very good job, but then starting from index of 0, you get 16 terms 0,1,2,3,4,5 ...15

OpenStudy (solomonzelman):

I would put\[\sum_{n=1}^{15}~~-9+5(n-1)\] I like index of 1 better:) They make more sense to me.

OpenStudy (anonymous):

but that isn't an answer choice >.<

OpenStudy (anonymous):

thanks by the way, can you help me with a few more? @SolomonZelman

OpenStudy (solomonzelman):

Yes, perhaps I can. I am not very good at math. Shoot ...!

OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence. Show all work for full credit. 1, -4, -9, -14, . . .

OpenStudy (solomonzelman):

pattern is? (you tell me)

OpenStudy (anonymous):

-5

OpenStudy (solomonzelman):

yes obviously you got this. Now, find the 12th term for me please.

OpenStudy (anonymous):

is there a short cut to finding it??

OpenStudy (anonymous):

if so, i don't know it TT^TT

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm a_n=a_1 +d(n-1) }\)

OpenStudy (solomonzelman):

you saw this before I am sure...

OpenStudy (anonymous):

ohhhhhhhhh yes that is very familiar

OpenStudy (solomonzelman):

yes, so for 12th term, what would you do, using the \(\large\color{purple}{ \rm a_n=a_1 +d(n-1) }\)

OpenStudy (anonymous):

\[a_{12}=1+5(12-1)\]??

OpenStudy (anonymous):

12th term= 56??

OpenStudy (solomonzelman):

yes !!!!

OpenStudy (anonymous):

yay ^_^

OpenStudy (solomonzelman):

Now, another formula.... \(\large\color{midnightblue}{ \rm S_n=\color{red}{\frac{1}{2}(a_1+a_n)}\times n }\)

OpenStudy (solomonzelman):

Just the concept, it would make to say that the red part is the average term (in arithmetic sequence) right ?

OpenStudy (solomonzelman):

it would make sense to say that the red, is the average term ?

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

So, can you find the sum ?

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm S_n=\color{red}{\frac{1}{2}(a_1+a_n)}\times n }\)

OpenStudy (solomonzelman):

Hold on !

OpenStudy (solomonzelman):

you found the 12th term incorrectly. you subtract 5 to get to the next term, a12 = a1 + (-5)(12-1) = a1 + (-55) = a1 - 55 = .... = -54

OpenStudy (anonymous):

OH!!!! ok!!

OpenStudy (solomonzelman):

now you can use the formula, \(\large\color{midnightblue}{ \rm S_n=\color{red}{\frac{1}{2}(a_1+a_n)}\times n }\) \(\large\color{midnightblue}{ \rm S_{12}=\color{red}{\frac{1}{2}(1+(-54)~~)}\times 12 }\)

OpenStudy (anonymous):

\[S_{12}=-318\]??

OpenStudy (solomonzelman):

YES !

OpenStudy (anonymous):

OH MY GOSH THANK YOU!!!! i love you so much >.<

OpenStudy (anonymous):

can you help me with one more??...Please ??

OpenStudy (solomonzelman):

Thank you, it's a pleasure to work with a smart person like you who can do the calculations (unlike so many other users on this website).

OpenStudy (solomonzelman):

Yes, I think I can.

OpenStudy (anonymous):

haha thanks >_< !!

OpenStudy (anonymous):

A certain radioactive isotope has a half-life of 11 days. If one is to make a table showing the half-life decay of a sample of this isotope from 32 grams to 1 gram; list the time (in days, starting with t = 0) in the first column and the mass remaining (in grams) in the second column, which type of sequence is used in the first column and which type of sequence is used in the second column?

OpenStudy (anonymous):

sometimes i just need a boost in the right direction and nudges on the way XP

OpenStudy (solomonzelman):

Ohh, I am sorry, I am not messing with science even if related to math. I used to be good at doing half lives, but I can't help you with this. Esp not with isotopes. very sorry • • • • •

OpenStudy (anonymous):

its ok, do you know anyone who might be able to help?? and thank you soooooooooo much for your help!!

OpenStudy (solomonzelman):

Not really, well you can look back into mathematics, and try to find someone of the top 99ers. I keep seeing people like me that do all maths but don't play games with science-:( Good luck !

OpenStudy (anonymous):

Thanks :D